# Variance-reduction for Variational Inequality Problems with Bregman Distance Function

### v2

@@ 4.1 Monotone VI
$\langle\bar{r}_{K},x_{K}-x\rangle\leq\|r_{K}\|_{\mathcal{X}^{*}}\|x-x_{K}\|_{\mathcal{X}}\leq{(1-\beta)L}\mathbf{D}_{X}(x_{K},x_{K-1})+\beta LD_{X}^{K-1}(x_{K})+L\mathbf{D}_{X}(x,x_{K}),$ (25)
@@
where the last inequality follows from unbiasedness of $\delta_{k}$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r}_{k}-\bar{r}_{k}$ and dropping the non-positive terms. Therefore, combining (27) with (26) and taking supremum over $x\in\tilde{X}_{R}$ followed by taking the expectation of both sides of the aforementioned inequality imply that
@@ 4.2 Non-monotone VIs
Therefore, from Assumption 4 followed by applying Lemma 1 parts (b) and (d) we conclude that
@@
$\leq(1-\beta)qL\mathbb{E}[\mathbf{D}_{X}(x_{k+1},x_{k})]+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4(1-\beta)L}}\sum_{t=i_{k}q+1}^{k}\mathbb{E}[\mathbf{D}_{X}(x_{t},x_{t-1})]+\theta_{k}\mathbb{E}[\left\langle r_{k},x_{k}-x_{k+1}\right\rangle]$
$\leq(\theta_{k}+{(1-\beta)q})L\mathbb{E}[\mathbf{D}_{X}(x_{k+1},x_{k})]+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4(1-\beta)L}}\sum_{t=i_{k}q+1}^{k}\mathbb{E}[\mathbf{D}_{X}(x_{t},x_{t-1})]$
@@
where $\mathcal{C}^{k}_{i}$’s are defined in Lemma 7.
@@ 5.1 Distributionally Robust Optimization
Table 3: Different datasets used in the experiment from LIBSVM [9].
@@ 5.2 Matrix Game
A first-order stationary solution of the above problem can be found by solving the corresponding VI formulated as (1) by letting $x=\begin{bmatrix}u\\ w\end{bmatrix}$, $F(x)=\begin{bmatrix}A^{\top}w-\upsilon u\\ -Au-\upsilon w\end{bmatrix}$, $h\equiv 0$, and $X=\{u\in\mathbb{R}^{n}\mid\left\|u\right\|_{2}\leq 1\}\times\{w\in\mathbb{R}^{m}\mid\left\|w\right\|_{2}\leq 1\}$. It is easy to verify that $F$ is a non-monotone and Lipschitz continuous map. In particular, similar to [36] we can show that the Lipschitz constant of $F$ is $L=\sqrt{\upsilon^{2}+\left\|A\right\|_{2}^{2}}$ and $F$ satisfies Assumption 4 with constant $\rho=\frac{\upsilon}{\upsilon^{2}+\left\|A\right\|_{2}^{2}}$. In our experiment, we let $\upsilon=1$, generate $A$ randomly from a standard Gaussian distribution then normalized such that $\left\|A\right\|_{2}=40$, $n=m$, and $F_{i}(x)=\begin{bmatrix}nA_{i:}w_{i}-\upsilon u\\ -mA_{:i}u_{i}-\upsilon w\end{bmatrix}$. We compare our proposed method with VR-MP and BC-SEG+ [35]. The parameters of our method are selected based on Remark 4, i.e., $\gamma=\beta=1-\frac{1}{2\sqrt{n}+1}$ and $\sigma=\frac{\nu}{6L}$ for some $\nu>0$. For VR-MP we select the step size as $\sigma=\frac{\nu}{\sqrt{n}L}$ and for BC-SEG+ we used $\alpha_{k}=\nu/k$. In our experiment, we select $\nu\in\{0.1,1,10\}$ as a tuning parameter to observe the performance of the considered methods with different step-sizes. We tested the performance of the considered methods for various dimensions $n=m\in\{100,500,1000\}$ and the plots are depicted in Figure 2. Since $x=\mathbf{0}$ is the equilibrium of this problem, the plots are in terms of the normalized solution norm, i.e., $\left\|x_{k}\right\|/\sqrt{n}$, versus the number of sample operators evaluated by the algorithms. Conversely, BC-SEG+ demonstrates convergence in lower dimensions but exhibits sluggish convergence when $n=1000$, possibly due to increased variance in sampling and a diminishing step-size. Our proposed method, VR-FoRMAB, consistently converges to the solution across all experiments, surpassing other methods.
@@ 6 Conclusions
In this paper, we studied VI problems with a finite-sum structure. Motivated by the lack of a unified method capable of adapting a Bregman distance function to address both monotone and non-monotone settings, we introduced a novel stochastic method called Variance-Reduced Forward Reflected Moving Average Backward (VR-FoRMAB). We established the convergence rate and oracle complexity of our method for the monotone case, achieving the optimal complexity of $\mathcal{O}(\sqrt{n}/\epsilon)$. Additionally, we analyzed the convergence of our algorithm in the non-monotone setting under the weak Minty Variational Inequality (MVI) assumption and the Lipschitz continuity of the Bregman distance generating function. Our proposed method attains a complexity of $\mathcal{O}(n/\epsilon^{2})$, offering an improvement in oracle complexity over existing approaches.

### v3

@@ 4.1 Monotone VI
$\langle\bar{r}_{K},x_{K}-x\rangle\leq\|\bar{r}_{K}\|_{\mathcal{X}^{*}}\|x-x_{K}\|_{\mathcal{X}}\leq{(1-\beta)L}\mathbf{D}_{X}(x_{K},x_{K-1})+\beta LD_{X}^{K-1}(x_{K})+L\mathbf{D}_{X}(x,x_{K}),$ (25)
@@
where the last inequality follows from unbiasedness of $\delta_{k}$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r}_{k}-\bar{r}_{k}$ and dropping the non-positive terms. Therefore, combining (26) with (27) and taking the supremum over $x\in\tilde{X}_{R}$ followed by taking the expectation of both sides of the aforementioned inequality implies that
@@ 4.2 Non-monotone VI
The recent studies on the weak Minty condition have shown that it holds under several structured conditions. In particular, this condition is satisfied when the operator $F$ is Lipschitz continuous and exhibits properties such as negative comonotonicity or positive cohypomonotonicity [6, 9]. Furthermore, in the context of min-max optimization problem, the weak Minty condition (with $\rho=0$) contains all quasiconvex-concave and starconvex-concave problems [18]. For further examples and applications see [12, 15, 41].
@@
Therefore, from Assumption 4 followed by applying Lemma 1 parts (a) and (d) we conclude that
@@
$\leq(1-\beta)qL\mathbb{E}[\mathbf{D}_{X}(x_{k+1},x_{k})]+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4(1-\beta)L}}\sum_{t=i_{k}q+1}^{k}\mathbb{E}[\mathbf{D}_{X}(x_{t},x_{t-1})]+\theta_{k}\mathbb{E}[\left\langle r_{k},x_{k}-x_{k+1}\right\rangle].$
The last term on the right-hand side of the above inequality can be bounded using Young’s inequality. Note that $r_{k}$ is either $F_{S}(x_{k})-(1-\beta)F_{S}({x}_{k-1})-\beta F_{S}(\tilde{x}_{k-1})$ or $\bar{r}_{k}$ which for both cases the inner product can be upper bounded similarly using Lipschitz continuity Assumption 1 as follows:
$\mathbb{E}[\left\langle r_{k},x_{k}-x_{k+1}\right\rangle]\leq\frac{(1-\beta)L}{2}\mathbb{E}[\|x_{k}-x_{k-1}\|_{\mathcal{X}}^{2}]+\frac{\beta L}{2}\mathbb{E}[\|x_{k}-\tilde{x}_{k-1}\|_{\mathcal{X}}^{2}]+\frac{L}{2}\mathbb{E}[\|x_{k+1}-x_{k}\|_{\mathcal{X}}^{2}].$ (34)
Therefore, combining (33) and (34), using strong convexity of Bregman distance function and Lemma 2 we conclude that
((*)( ∗ ) $\leq(\theta_{k}+{(1-\beta)q})L\mathbb{E}[\mathbf{D}_{X}(x_{k+1},x_{k})]+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4(1-\beta)L}}\sum_{t=i_{k}q+1}^{k}\mathbb{E}[\mathbf{D}_{X}(x_{t},x_{t-1})]$ (35)
@@
where $\mathcal{C}^{k}_{i}$’s are defined in the statement of Lemma 7.
@@ 5.1 Distributionally Robust Optimization
Table 3: Different datasets used in the experiment from LIBSVM [8].
@@ 5.2 Matrix Game
A first-order stationary solution of the above problem can be found by solving the corresponding VI formulated as (1) by letting $x=\begin{bmatrix}u\\ w\end{bmatrix}$, $F(x)=\begin{bmatrix}A^{\top}w-\upsilon u\\ -Au-\upsilon w\end{bmatrix}$, $h\equiv 0$, and $X=\{u\in\mathbb{R}^{n}\mid\left\|u\right\|_{2}\leq 1\}\times\{w\in\mathbb{R}^{m}\mid\left\|w\right\|_{2}\leq 1\}$. It is easy to verify that $F$ is non-monotone, since the symmetric part of its Jacobian is indefinite when $A$ is non-symmetric. Also, similar to  [41], we can show $F$ is Lipschitz continuous with constant $L=\sqrt{\upsilon^{2}+\|A\|_{2}^{2}}$, and $F$ satisfies Assumption 4 with constant $\rho=\frac{\upsilon}{\upsilon^{2}+\left\|A\right\|_{2}^{2}}$. In our experiment, we let $\upsilon=1$, generate $A$ randomly from a standard Gaussian distribution then normalized such that $\left\|A\right\|_{2}=40$, $n=m$, and $F_{i}(x)=\begin{bmatrix}nA_{i:}w_{i}-\upsilon u\\ -mA_{:i}u_{i}-\upsilon w\end{bmatrix}$. We compare our proposed method with VR-MP and BC-SEG+ [40]. The parameters of our method are selected based on Remark 4, i.e., $\gamma=\beta=1-\frac{1}{2n+1}$ and $\sigma=\frac{\nu}{6L}$ for some $\nu>0$. For VR-MP we select the step size as $\sigma=\frac{\nu}{L}$ and for BC-SEG+ we used $\alpha_{k}=\nu/k$. In our experiment, we select $\nu\in\{0.1,1,10\}$ as a tuning parameter to observe the performance of the considered methods with different step-sizes. We tested the performance of the considered methods for various dimensions $n=m\in\{100,500,1000\}$ and the plots are depicted in Figure 2. Since $x=\mathbf{0}$ is the equilibrium of this problem, the plots are in terms of the normalized solution norm, i.e., $\left\|x_{k}\right\|/\sqrt{n}$, versus the number of sample operators evaluated by the algorithms. Conversely, BC-SEG+ demonstrates convergence in lower dimensions but exhibits sluggish convergence when $n=1000$, possibly due to increased variance in sampling and a diminishing step-size. Our proposed method, VR-FoRMAB, consistently converges to the solution across all experiments, surpassing other methods.
@@ 6 Conclusions
In this paper, we studied VI problems with a finite-sum structure. Motivated by the lack of a unified method capable of adapting a Bregman distance function to address both monotone and non-monotone settings, we introduced a novel stochastic method called Variance-Reduced Forward Reflected Moving Average Backward (VR-FoRMAB). We established the convergence rate and oracle complexity of our method for the monotone case, achieving the optimal complexity of $\mathcal{O}(n+\frac{\sqrt{n}}{\epsilon})$. Additionally, we analyzed the convergence of our algorithm in the non-monotone setting under the weak Minty Variational Inequality (MVI) assumption and the Lipschitz continuity of the Bregman distance generating function. Our proposed method attains a complexity of $\mathcal{O}(n+\frac{1}{\epsilon^{2}})$, offering an improvement in oracle complexity over existing approaches.
Funding
This work was supported by the National Science Foundation (NSF) under Grant ECCS-2231863.
Disclosure statement
The authors report there are no competing interests to declare.